OpenStudy (anonymous):

simplify to the nearest ten-thousandth: log4 (2961)

OpenStudy (anonymous):

I think the equation you need is: \[\log_{a} x = \ln x / \ln a\] So, we can convert your problem into the following then evaluate it: \[\log_4 2961 = \ln 2961 / \ln 4\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Your'e welcome, glad I could help

OpenStudy (reemii):

calculator allowed?